The beat setup has been made on the PSL table. The BS and the PD were setup. The beat was found at 29.42degC and 50.58degC for the PSL and AUX crystal temperatures, respectively.
We are ready for the EOM test. I have instruments stacked around the PSL table. Please leave them as they are for a while. If you need to move them, please contact with me. Thanks.
A picked-off PSL after the main modulator was used as the PSL beam. This was already introduced close to the setup thanks to the previous 3f cancellation test ELOG 11029. The AUX beam was obtained from the transmission of 90% mirror. Both paths have S polarization. The beams are combined with a S-pol 50% BS. The combined beam is detected by a new focus 1GHz PD.
The PSL crystal temp (actual) was 50.58degC. The AUX crystal temp was swept upward and the string beat was found at 50.58degC. After a bit of alignment, the beat strength was -18dBm (at 700V/A RF transimpedance of NF1611) .